A violin plays a note at 70 db, then a second, third, and fourth violin join in succession to play the same one each at 70 db. How much louder in decibels does the sound get as each violin joins in, and how much louder is the total sound?
Answer #1

Sound level in dB is given by

B=10log(I/Io)

70=10log(I/10-12)

I=10-12*107=1*10-5W/m2

Total intensity

I=10-5*4 =4*10-5 W/m2

Therefore total sound intensity is

Btotal=10Log(4*10-5/10-12)=76.02 dB

And total sound is louder by

dB =76.02-70=6.02 dB